4. Valle de Vallibierna

’ANETO’ PEAK 5 hours 30 mins. 3,404 m. high

The Aneto is the highest peak of the Pyrenees. To reach it, it is necessary to cross a small glacier and therefore you will need to take with you all necessary high mountain tools such as ropes, ice axe and crampons.

The departure point is the end of the asphalt trail that goes to the Ospital, Besurta and the Rencllusa. Take the track on your right to the refuge of the Rencllusa (2,140 m.) which you will reach in about 30 minutes.

Keep going up towards the crest of the Portillons on your left side and you will reach the Lower Portillón (2,745 m.) in 2 hours and 15 minutes. Keep following the right (West) side of the crest till you get to the Upper Portillón (2,870 m.) after 3 hours. Cross the Portillón towards the glacier of the Aneto, which you will cross diagonally to the Collada de Coronas (3,198 m. - 4 h.45 min.). From there, go up the left (North) side of the crest until, virtually on its top, it narrows into the well-know Mahoma Bridge, which has to be crossed very carefully. Immediately after having crossed it you will reach the top of the Aneto.

High Mountain Excursion. It is essential that you be accompanied by a High Mountain Guide (if you do not have previous experience), wear boots and material such as crampons and ice axe. Not recommended for children.